In order to build a facility you take out a $100mil loan 15% interest and issues $50mil worth of bonds for which you have to pay 10% interest. What is your WACC (weighted-average cost of capital)?
Answer #1
Total Capital value = Value of Loan + Value of Debt
=100+50
=150
Weight of Loan = Value of Loan/Total Capital Value
= 100/150
=0.6667
Weight of Debt = Value of Debt/Total Capital Value
= 50/150
=0.3333
Cost of Capital = Weight of Loan*Cost of Loan+Weight of Debt*Cost of Debt
Cost of Capital = 15*0.6667+10*0.3333
Cost of Capital = 13.3333